CAT Chat Minutes

January 23, 1998

Issues that require action and/or follow-up

1. The central LOM door has been oscillating back and forth. What is causing this problem?

We will call and have the problem checked. We have called for maintenance whenever a problem is brought up at CAT Chat and will continue to do so. In 435, we replaced the sensor with a less sensitive one. The company will be out on Friday to begin correcting the door problems.

2. Fire extinguishers are being installed on the interior of the columns; some of these are being put up on the only wall space that we have. Could these be installed in areas that are already "keep clear" areas?

The "keep clear" area is where they are being installed. UNI Cat may be the exception to this. Some electrical work must be done in that sector. When that work is complete, the fire extinguisher will be relocated.

3. The x-ray BPMs have been unreliable for the past month. They have given readouts that are unreliable. When will these be more reliable?

In Sector 17, we have done so much orbit correction that the x-ray beam has moved towards one set of BPM blades, making the BPM inoperable. We will correct this by moving the x-ray BPM, but we need time from Sector 17 to do this. We will work with Sector 17 personnel to correct this problem.

4. Sector 9 asked if everyone went to the high power undulator, would we have enough rf power for everyone?

Yes, we would have enough power. When we developed the plan for the APS, we designed for more wigglers, which need higher rf power, in addition to the 5-m-long IDs. Hence there is adequate rf power available.

5. Since we have more CATs coming on board, will there be construction of more LOMs?

Yes, we have an LOM planned that will support an additional 4 sectors. The new LOM (436) will support sectors 21 - 24.

6. Is there a time table for the installation of the liquid nitrogen systems yet?

Bill Osterhuis, Program Manager at DOE, has informed us that the funding will become available over two years. The first installment may become available in the later part of 1998.

7. While we are waiting for the partitions that will be put on the floor (to separate sectors from the main aisle), can we put our storage cabinets up to the area that will be occupied by the partitions?

Yes, you may do this. The cabinets must be under 7 ft and must open to the inside.

Information/Follow-up Items

APS Operational Status Update:  We will continue the run as scheduled. Monday at 0800 hrs, we are scheduled for 24 hours of Machine Studies. Tuesday at 0800, we have shielding verification at 25 mA for one shift, then back to normal.

General:  Over the weekend (Jan. 17, 1998) Commonwealth Edison had an explosion in one of their power plants. DuPage County experienced a significant but short term drop in power that caused many systems at the APS to trip off. These situations are rare with ComEd. We managed to get the entire facility running within six hours.

One of the new rf systems is having some harmonic problems. We will attempt a repair during the Machine Studies on Monday. The 3-ID beamline is most affected, because the problem only affects bunch purity.

The gateway problems have continued. We found a problem with how the gateway starts after reboot and problems with some PVs, some of which could be accessed and some that could not be. The current configuration requires that all controlling PVs have ".VAL" field for access. This has been fixed. Once you make a request for a certain PV, the gateway makes the connection and holds it for 20 minutes. We are still noticing loss of connection when an IOC having access security is rebooted. Work is in progress to resolve this problem.

Everyone should be on the new gateway, which we guarantee not to touch during a run.

We will continue to work on the direct link for ID gap control. It is still too early to give you a hard date as to when it will be ready.

We are still experiencing some stalls in the undulator gap motion due to software glitches and hardware issues. During the down time, we did maintenance on many of the devices. This is the reason in part for the stalls. We thought we had it debugged, but, as we are running, we are finding we still have some problems.

If you have any further input on the ID beamline energy calibration, please contact M. Ramanathan by e-mail ( or talk with your FC.

We have the User Steering Committee meeting on Thursday, Jan. 29, 1998, and the Research Directorate meeting on Friday, Jan. 30, 1998. If you have any issues, please convey them to your CAT Directors.